OpenStudy (anonymous):

Each side of a square loop of wire measures 2.0 cm. A magnetic field of 0.044 T perpendicular to the loop changes to zero in 0.10 s. What average emf is induced in the coil during this change? A. 1.8 V B. 0.088 V C. 0.88 V D. 0.00018 V

OpenStudy (michele_laino):

the magnetic flux change is: \[\Large \Delta \Phi = S \times \Delta B = {\left( {2 \times {{10}^{ - 2}}} \right)^2} \times 0.044 = ...Weber\]

OpenStudy (michele_laino):

\[\large \Delta \Phi = S \times \Delta B = {\left( {2 \times {{10}^{ - 2}}} \right)^2} \times 0.044 = ...Weber\]

OpenStudy (anonymous):

1.76E-5? choice D is the solution?

OpenStudy (michele_laino):

no, since we have to find the emf

OpenStudy (michele_laino):

more explanation:

OpenStudy (anonymous):

ohh ok! how do we do taht?

OpenStudy (michele_laino):

before magnetic flux chnaging, the flux of the magnetic field through the square loop is: area*magnetic field=0.02*0.02*0.044 after the magnetic field changing the new magnetic flux through the square loop is: area*magnetic field=0.02*0.02*0=0 

OpenStudy (michele_laino):

so the requested emf is: \[\Large E = \frac{{\Delta \Phi }}{{\Delta t}} = ...volts\] that is the Faraday-Neumann law

OpenStudy (anonymous):

oh ok! what do we plug in? :/

rvc (rvc):

emf= flux/area

OpenStudy (michele_laino):

no, emf= flux/time

rvc (rvc):

oh ye

OpenStudy (michele_laino):

please apply the Faraday-Neumann law @rvc

OpenStudy (anonymous):

what are we plugging in? :/

rvc (rvc):

yes faradays law applies here :)

rvc (rvc):

i just messed the equation :)

OpenStudy (michele_laino):

next step, is: \[\Large E = \frac{{\Delta \Phi }}{{\Delta t}} = \frac{{0.176 \times {{10}^{ - 4}}}}{{0.1}} = ...volts\]

OpenStudy (michele_laino):

no worries!! :) @rvc

OpenStudy (anonymous):

1.76 E-4? so the solution is choice D?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!:)

rvc (rvc):

you are the BEST @Michele_Laino

OpenStudy (michele_laino):

:) @rvc